# Comparing sampling techniques to chart parameter space of 21 cm Global signal with Artificial Neural Networks

### v1

@@ 1 Introduction
Over recent years, machine learning (ML) algorithms have gained extensive popularity in signal modelling and parameter estimation. For signal modelling, approaches such as those by [20, 21, 22, 23, 24] leverage Artificial Neural Networks (ANN) to model the 21cm signal across various aspects. Simultaneously, [25, 26, 27] employ ML algorithms to extract parameters linked to the signal. Apart from the ML algorithm, several traditional methods are used to remove the signal and associated parameters [28, 29, 30, 31]. Currently, due to limited observational constraints on intergalactic medium (IGM) properties, there is no singular and firmly established set of quantities to parameterize the underlying astrophysical processes shaping the signal. Several potential parameters have been suggested in various proposed models of the signal[32, 33, 34, 35].
@@ 2 HI 21cm Signal
The spin temperature, which is the relative populations of hydrogen atoms in the two spin states, is decided by competition between 3 processes and their corresponding physical quantities: (1) absorption of CMB photons and stimulated emission; $T_{CMB}$, the temperature of CMB, (2) collisions with other hydrogen atoms (H-H), free electrons (H-e), free hydrogen nuclei (H-p); $T_{k}$, Kinetic gas temperature (of IGM), and (3) scattering of Lyman- $\alpha$ photons, $T_{\alpha}$, colour temperature for Wouthuysen–Field effect. The spin temperature, $T_{s}$ is calculated as [37][38]:
@@ 4 Models for the Global 21cm Signal
We have used the tanh model to simulate the global 21 cm signal during the cosmic dawn (CD) and the EoR using ARES [41]. The model uses simple parametric forms for the Lyman $\alpha$ background, IGM temperature, and re-ionization histories [42]. This method models the signal using IGM properties like the strength of Lyman $\alpha$ coupling, $J_{\alpha}(z)$; the temperature of the IGM, $T(z)$; and ionization fraction, ${\overline{X}_{i}}$, but do not take into account the source properties. Each of these quantities are evolved as:
@@
The step height, $A_{ref}$, in units of $10^{-21}$ $\rm ergs^{-1}cm^{-2}Hz^{-1}sr^{-1}$ corresponding to the Lyman $\alpha$ ($Ly\alpha$) background is $J_{ref}$. It saturates at low redshift, with a value of 11.69 as determined by [42] using MCMC parameter estimation. In our study, we treat it as a constant, as done in our work. The redshift interval and pivot redshift for $Ly\alpha$ background tanh parametrization are $J_{dz}$ and $J_{z0}$ respectively. Similarly, for X-ray heating, the temperature of IGM, $T(z)$, $T_{dz}$ and $T_{z0}$ denotes redshift interval and central redshift in units of Kelvin. The amplitude, $A_{ref}$, corresponding to IGM temperature, represented by $T_{ref}$, saturates at around 1000K. For the ionization fraction, ${\overline{X}_{i}}$, the natural value for step height is unity. The redshift interval $\Delta z$ and the pivot redshift $z_{0}$ over which ionization takes place are given by $X_{dz}$ and $X_{z0}$, respectively. The values for these parameters inferred by [42] are: $J_{dz}$ = 3.31, $J_{z0}$ = 18.54, $T_{dz}$ = 2.82, $T_{z0}$ = 9.77, $X_{dz}$ = 2.83, $X_{z0}$ = 8.68.
@@ 4.2 Astrophysical Model
    Figure 1: Each subplot illustrates training datasets for the global 21cm signal, constructed using Random, Latin Hypercube, and Hammersley sequence sampling methods, respectively, for the parameterized signal. The final subplot depicts training datasets with added foreground and thermal noise. Signal subsets are highlighted in color, while the remaining sets are displayed in gray as the background.
@@ 6 Artificial Neural Networks
Artificial Neural Networks (ANNs) are widely used supervised machine learning algorithms, with applications increasingly found in cosmology as well. Mathematically, for a given set of [$x_{i},y_{i}$], ANNs try to find a function, f(x), such that $y_{i}=f(x_{i})$ through a series of weighted summations. This operation occurs in a series of layers. The neurons are the building blocks of the ANN. The ANN contains n number of layers, with one input layer, one output layer, and n-2 hidden layers. The input values $\mathbf{x_{i}}$ with D dimensions are given to the first layer with D nodes. The output from one layer $a_{i}$ goes to the next layer’s $j^{th}$ neuron as input $a_{j}$:
$a_{j}=h\left(\sum_{i=1}^{D}w_{ji}x_{i}\right)+w_{j0}$ (6.1)
Here, $h$ is the activation function, $w_{ji}$ and $w_{j0}$ are weights and biases associated with the neuron set, respectively. There is no activation function used in the output layer.
@@
$\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i}^{(n)})^{2}$ (6.2)
Here, $y_{i}$ is the true output label and $\hat{y}_{i}^{(n)}$ is the prediction after updating the weights, $w_{ji}$, and biases, $w_{j0}$, with n iterations of the ANNs:
$\hat{y}_{i}^{(n)}=f\left(\sum_{j=1}^{m}w_{ji}^{(n)}\cdot x_{i}+w_{j0}^{(n)}\right)$ (6.3)
Most optimizers are based on the gradient descent method of loss optimization.
The number of layers, number of neurons in each layer, learning rate, and activation function are to predecided. These parameters are known as hyperparameters. We can search over a range of hyperparameters and use the ones giving minimum loss and maximum accuracy. In this work, Scikit-learn  [^2] https://scikit-learn.org/stable/ and Keras [^3] https://keras.io/ are used for building an ANN.
@@ 6.1 Metric of accuracy
For testing and validating the data set, RMSE and R squared value is used as a measure to see the difference between predictions of input parameters made by the model for testing data and the actual value.
@@ 7 Training and Testing data set
Table 1: The range of parameters used to build the training dataset for the Parametrized and Non-parameterized (Physical) cases of global 21cm signals and foregrounds.
@@ 8.3 Generalizability Test of the ANN Models
Similarly, for the parameterized signal and foreground case, ANN models trained with Hammersley sampled datasets consistently exhibited high prediction accuracy across all test datasets, with the highest $\rm R^{2}$ score approximately around 0.92 and the lowest around 0.88, with a mean of approximately 0.91 for each case. Detailed results are depicted in Fig. 9 in the second row. Conversely, ANN models trained with Latin Hypercube and Random sampled datasets showed inconsistent prediction accuracy for both cases. For test datasets generated with HSS, the lowest $\rm R^{2}$ score was less than 0.70, approximately 19-20 % lower than the mean $\rm R^{2}$ score, approximately 0.90. Similar fluctuations were observed for test datasets constructed using Latin Hypercube and Random sampling methods; detailed results are presented in Fig. 9, with the colour scheme remaining consistent across all subplots. Similarly, for the non-parameterized signal case, we observed consistent behaviour, with ANN models trained on Hammersley sampled datasets exhibiting less deviation from the mean $\rm R^{2}$ score compared to models trained on LHS and Random datasets, the detailed result present in Appendix A.3.
@@ Acknowledgements
AT would like to thank the Indian Institute of Technology Indore for providing funding for this study in the form of a Teaching Assistantship. SM and AD acknowledge financial support through the project titled “Observing the Cosmic Dawn in Multicolour using Next Generation Telescopes” funded by the Science and Engineering Research Board (SERB), Department of Science and Technology, Government of India through the Core Research Grant No. CRG/2021/004025.
@@ A.2 Signal with Foreground and Thermal Noise
Similar to the parameterized case, we introduced the effects of foreground and thermal noise in the non-parametric signal. Given fewer free parameters than the parametric case, we generated datasets in three sizes: 10,000, 50,000, and 100,000 samples. The architecture of the ANN model closely resembles that of the parameterized case, with the primary difference being the output layer of the second ANN model, which comprises 5 neurons. Each neuron represents 3 signal parameters and 2 foreground parameters ($a_{0}$, $a_{3}$). Here, we observe a consistent trend similar to parametric scenarios: the ANN model trained with 10000 datasets achieves an overall $\rm R^{2}$ score of around 0.92, which increases to approximately 0.94 with 50000 datasets, and around 0.96 with 10,$0$ datasets. The detailed results for each sampling method with the various dataset sizes, the $\rm R^{2}$ and RMSE score for individual parameters, are listed in Tab. 8 and Tab. 9. We also individually visualized the predicted parameter values against the original values for each sampling method across different dataset sizes. These visualizations are presented in Fig. 13 for HSS, Fig. 14 for LHS, and Fig. 15 for Random Sampling.
@@ A.3 Generalizability Test
We have conducted a generalizability test to demonstrate the robustness of the ANN model for non-parameterized scenarios. The detailed results are plotted below in the histogram, see Fig. 16.
Figure 16: The figure illustrates ANN model predictions for various trials trained on datasets sampled using HSS, LHS, and Random methods with optimal sample sizes. Histograms depict ANN prediction accuracy measured by $\rm R^{2}$ scores. Panels in the figure’s rows correspond to different scenarios: the top three panels in the first row show predictions for the Non-parameterized signal, followed by predictions for the Non-parameterized signal with foreground and thermal noise in the second row. Histograms are colour-coded: blue represents HSS-trained ANN models, orange represents LHS-trained ANN models, and green represents the Random-trained ANN model’s prediction accuracy.
@@ Appendix B Training Loss and Validation Loss of the ANN models
Figure 17: The figure illustrates the evolution of the network’s loss function across various scenarios. The training loss is denoted by a solid line, and the validation loss is indicated by a dashed line over epochs. Notably, the test loss closely follows the training loss in this visualization. Top row: training loss for parameterized signal (left) and non-parameterized signal (right). Second row: training and validation loss for parameterized signals with foreground. Bottom row: training and validation loss for foreground-added non-parameterized signals.

### v2

@@ 1 Introduction
Over recent years, machine learning (ML) algorithms have gained extensive popularity in signal modelling and parameter estimation. For signal modelling, approaches such as those by [18, 19, 20, 21, 22, 23] leverage Artificial Neural Networks (ANN) to model the 21cm signal across various aspects. Simultaneously, [24, 25, 26, 27, 28] employ ANN algorithms to extract parameters linked to the signal. Additionally, simulation-based inference (SBI) techniques, leveraging deep learning, are increasingly preferred for astrophysical inference from cosmic 21 cm signals, particularly for retrieving posteriors of astrophysical parameters via likelihood-free Bayesian inference directly from power spectrum and light-cone image analysis [29, 30, 31, 32, 33]. [34] introduced the Solid Harmonic Wavelet Scattering Transform (WST) and ”3D ScatterNet” to enhance the inference of astrophysical parameters. Apart from the ML algorithm, several traditional methods are used to remove the signal and associated parameters [35, 36, 37, 38]. Currently, due to limited observational constraints on intergalactic medium (IGM) properties, there is no singular and firmly established set of quantities to parameterize the underlying astrophysical processes shaping the signal. Several potential parameters have been suggested in various proposed models of the signal[39, 40, 41, 42].
@@ 2 HI 21cm Signal
The spin temperature, which is the relative populations of hydrogen atoms in the two spin states, is decided by competition between 3 processes and their corresponding physical quantities: (1) absorption of CMB photons and stimulated emission; $T_{CMB}$, the temperature of CMB, (2) collisions with other hydrogen atoms (H-H), free electrons (H-e), free hydrogen nuclei (H-p); $T_{k}$, Kinetic gas temperature of IGM, and (3) scattering of Lyman- $\alpha$ photons, $T_{\alpha}$, colour temperature for Wouthuysen–Field effect. The spin temperature, $T_{s}$ is calculated as [6][43]:
@@ 4 Models for the Global 21cm Signal
We have used the tanh model to simulate the global 21 cm signal during the cosmic dawn (CD) and the EoR using ARES [46]. The model uses simple parametric forms for the Lyman $\alpha$ background, IGM temperature, and re-ionization histories [47]. This method models the signal using IGM properties like the strength of Lyman $\alpha$ coupling, $J_{\alpha}(z)$; the temperature of the IGM, $T(z)$; and ionization fraction, ${\overline{X}_{i}}$, but does not take into account the source properties. Each of these quantities is evolved as:
@@
The step height, $A_{ref}$, in units of $10^{-21}$ $\rm ergs^{-1}cm^{-2}Hz^{-1}sr^{-1}$ corresponding to the Lyman $\alpha$ ($Ly\alpha$) background is $J_{ref}$. It saturates at low redshift, with a value of 11.69, as determined by [47] using MCMC parameter estimation. In our study, we treat it as a constant, as done in our work. The redshift interval and pivot redshift for $Ly\alpha$ background tanh parametrization are $J_{dz}$ and $J_{z0}$ respectively. Similarly, for X-ray heating, the temperature of IGM, $T(z)$, $T_{dz}$ and $T_{z0}$ denotes redshift interval and central redshift in units of Kelvin. The amplitude, $A_{ref}$, corresponding to IGM temperature, represented by $T_{ref}$, saturates at around 1000K. For the ionization fraction, ${\overline{X}_{i}}$, the natural value for step height is unity. The redshift interval $\Delta z$ and the pivot redshift $z_{0}$ over which ionization takes place are given by $X_{dz}$ and $X_{z0}$, respectively. The values for these parameters inferred by [47] are: $J_{dz}$ = 3.31, $J_{z0}$ = 18.54, $T_{dz}$ = 2.82, $T_{z0}$ = 9.77, $X_{dz}$ = 2.83, $X_{z0}$ = 8.68.
@@ 4.2 Astrophysical Model
    Figure 1: Each subplot illustrates training datasets for the global 21cm signal, constructed using Random, Latin hypercube, and Hammersley sequence sampling methods, respectively, for the parametrized signal. The final subplot depicts training datasets with added foreground and thermal noise. Signal subsets are highlighted in color, while the remaining sets are displayed in gray as the background.
@@ 6 Artificial Neural Networks
Artificial neural networks (ANNs) are widely used supervised machine learning algorithms, with applications increasingly found in cosmology as well. Mathematically, for a given set of [$x_{i},y_{i}$], ANNs try to find a function, f(x), such that $y_{i}=f(x_{i})$ through a series of weighted summations. This operation occurs in a series of layers. The neurons are the building blocks of the ANN. The ANN contains n number of layers, with one input layer, one output layer, and n-2 hidden layers. The input values $x_{i}$ with D dimensions are given to the first layer with D nodes. The output from one layer $a_{i}$ goes to the next layer’s $j^{th}$ neuron as input $a_{j}$. The computation for the activation of the $j^{t}h$ neuron in layer $l+1$ is given by:
$a^{l+1}_{j}=h\left(\sum_{i=1}^{D}w_{ji}a^{l}_{i}+w_{j0}\right)$ (6.1)
Here, $h$ represents the activation function, $l$ denotes the layer index, and $a$ refers to the inputs for the corresponding layer. For l= 0, $a^{0}_{i}=x_{i}$ which represents the input values for the first layer. The terms $w_{ji}$ and $w_{j0}$ are weights and biases associated with the neuron set, respectively. There is no activation function used in the output layer.
@@
$\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}$ (6.2)
Here, n is number of samples, $y_{i}$ is the true output label and $\hat{y}_{i}$ is the prediction after updating the weights, $w_{ji}$, and biases, $w_{j0}$, with every iterations of the ANNs. The prediction $\hat{y}_{j}$ for the $j^{th}$ neuron in the output layer is calculated as:
$\hat{y}_{j}=\sum_{i=1}^{D}w_{ji}\cdot a^{L-1}_{i}+w_{j0}$ (6.3)
Here, L is the index of the final layer. Most optimizers are based on the gradient descent method of loss optimization.
The number of layers, number of neurons in each layer, learning rate, and activation function are to predecided. These parameters are known as hyperparameters. We can search over a range of hyperparameters and use the ones giving minimum loss and maximum accuracy. In this work, Scikit-learn  [^2] https://scikit-learn.org/stable/ and Keras [^3] https://keras.io/ are used for building an ANN. We optimized hyperparameters using trial and error to determine the best architecture for the ANNs. Details of the final architecture are provided in the following section.
@@ 6.1 Metric of accuracy
For testing and validating the data set, RMSE and R-squared value is used as a measure to see the difference between predictions of input parameters made by the model for testing data and the actual value.
@@ 7 Training and testing data set
Table 1: The range of parameters used to build the training dataset for the Parametrized and Non-parametrized (Physical) cases of global 21cm signals and foregrounds.
@@ 8.3 Generalizability test of the ANN models
Similarly, for the parametrized signal and foreground case, ANN models trained with Hammersley sampled datasets consistently exhibited high prediction accuracy across all test datasets, with the highest $\rm R^{2}$ score approximately around 0.92 and the lowest around 0.88, with a mean of approximately 0.91 for each case. Detailed results are depicted in Fig. 9 in the second row. Conversely, ANN models trained with Latin hypercube and Random sampled datasets showed inconsistent prediction accuracy for both cases. For test datasets generated with HSS, the lowest $\rm R^{2}$ score was less than 0.70, approximately 19-20 % lower than the mean $\rm R^{2}$ score, approximately 0.90. Similar fluctuations were observed for test datasets constructed using Latin hypercube and Random sampling methods; detailed results are presented in Fig. 9, with the colour scheme remaining consistent across all subplots. Similarly, for the non-parametrized signal case, we observed consistent behaviour, with ANN models trained on Hammersley sampled datasets exhibiting less deviation from the mean $\rm R^{2}$ score compared to models trained on LHS and Random datasets, the detailed result present in Appendix A.3.
@@ Acknowledgements
AT would like to thank the Indian Institute of Technology Indore for providing funding for this study in the form of a Teaching Assistantship. SM and AD acknowledge financial support through the project titled “Observing the Cosmic Dawn in Multicolour using Next Generation Telescopes” funded by the Science and Engineering Research Board (SERB), Department of Science and Technology, Government of India through the Core Research Grant No. CRG/2021/004025. The authors acknowledge the use of facilities procured through the funding via the Department of Science and Technology, Government of India sponsored DST-FIST grant no. SR/FST/PSII/2021/162 (C) awarded to the DAASE, IIT Indore.
@@ A.2 Signal with foreground and thermal noise
Similar to the parametrized case, we introduced the effects of foreground and thermal noise in the non-parametric signal. Given fewer free parameters than the parametric case, we generated datasets in three sizes: 10,000, 50,000, and 100,000 samples. The architecture of the ANN model closely resembles that of the parametrized case, with the primary difference being the output layer of the second ANN model, which comprises 5 neurons. Each neuron represents 3 signal parameters and 2 foreground parameters ($a_{0}$, $a_{3}$). Here, we observe a consistent trend similar to parametric scenarios: the ANN model trained with 10000 datasets achieves an overall $\rm R^{2}$ score of around 0.92, which increases to approximately 0.94 with 50000 datasets, and around 0.96 with 10,$0$ datasets. The detailed results for each sampling method with the various dataset sizes, the $\rm R^{2}$ and RMSE score for individual parameters, are listed in Tab. 8 and Tab. 9. We also individually visualized the predicted parameter values against the original values for each sampling method across different dataset sizes. These visualizations are presented in Fig. 13 for HSS, Fig. 14 for LHS, and Fig. 15 for Random sampling.
@@ A.3 Generalizability test
We have conducted a generalizability test to demonstrate the robustness of the ANN model for non-parametrized scenarios. The detailed results are plotted below in the histogram, see Fig. 16.
Figure 16: The figure illustrates ANN model predictions for various trials trained on datasets sampled using HSS, LHS, and Random methods with optimal sample sizes. Histograms depict ANN prediction accuracy measured by $\rm R^{2}$ scores. Panels in the figure’s rows correspond to different scenarios: the top three panels in the first row show predictions for the Non-parametrized signal, followed by predictions for the Non-parametrized signal with foreground and thermal noise in the second row. Histograms are colour-coded: blue represents HSS-trained ANN models, orange represents LHS-trained ANN models, and green represents the Random-trained ANN model’s prediction accuracy.
@@ Appendix B Training loss and validation loss of the ANN models
Figure 17: The figure illustrates the evolution of the network’s loss function across various scenarios. The training loss is denoted by a solid line, and the validation loss is indicated by a dashed line over epochs. Notably, the test loss closely follows the training loss in this visualization. Top row: training loss for parametrized signal (left) and non-parametrized signal (right). Second row: training and validation loss for parametrized signals with foreground. Bottom row: training and validation loss for foreground-added non-parametrized signals.